### v1

@@ 2.1 Source masking
We mask out all eRASS:4 X-ray sources in the 0.2–2.3 keV eSASS detection catalog [^1] Proprietary eROSITA data. Version all_s4_SourceCat1B_221031_poscorr_mpe_photom. The detection pipeline used in this work is the same as the one version used for the first release of the All-Sky Survey data (Merloni et al. 2024). The catalog contains point-like sources and extended sources, such as supernova remnants, galaxies, galaxy groups, and clusters. To be conservative, we do not apply selection cuts on the contaminating X-ray emitting sources.
We use a background level-dependent radius to exclude detected sources as follows. For each source, the surface brightness profile can be approximately described as
@@
where $r_{\mathrm{PSF}}=9^{\prime\prime}$ is the survey averaged point spread function (PSF) core radius parameter obtained from point source stacking and $r_{\mathrm{EXT}}$ is the parameter EXT given by the catalog if the source is extended. We adopt the radius at which the source profile reaches 0.1 (10%) of the background level as the masking radius, i.e., $S(r_{\mathrm{cut}})=0.1\times S_{\mathrm{bkg}}$, where $S_{\mathrm{bkg}}$ is obtained from ML_BKG_1 from the detection pipeline.
@@
For extended sources with large angular sizes, the eSASS detection pipeline usually returns several detections known as split sources (Liu et al. 2022, 2024). These large angular size sources are already included in the eRASS1 galaxy cluster and group catalog (Bulbul et al. 2024). Therefore, we mask out all sources in the eRASS1 galaxy cluster and group catalog with radii of $1.5\times r_{500}$. Additionally, we adopt an optical cluster catalog compiled by the redMaPPer algorithm (Rykoff et al. 2014) on the DESI Legacy Survey (Dey et al. 2019) DR9 + DR10 grz images in a blind mode (Kluge et al. 2024). We exclude sources with richness $\lambda>20$ and redshift $0.80.2<z<$ 0.8 in radii of $1.5\times r_{\lambda}$, where $r_{\lambda}=(\lambda/100)^{1/5}$ $h^{-1}$ Mpc, is the richness scaled radius from the redMaPPer algorithm. This cut safely removes clusters of galaxies with both high completeness and high purity, especially for high redshift systems whose X-ray luminosity is below the eRASS1 detection limit.
@@
To conservatively remove the emission from low-richness galaxy groups in the local Universe, we also adopt a galaxy group catalog (Tinker 2021) based on the SDSS Main Galaxy Sample (Strauss et al. 2002). We exclude sources with $n_{\mathrm{sat}}>1$ in radii of $1.5\times r_{500}$. If a source has $4001.5\times r_{500}<$ 400 kpc, we use a fixed value of 400 kpc instead. In addition to those cluster and group catalogs, we mask out critical points identified by DisPerSE of type maximum and bifurcation in the filament catalog in a redshift range $0.60.2<z<$ 0.6 with fixed radii of 2 Mpc as we consider these to be the nodes of the large scale structure (see Malavasi et al. 2020, for the details). The details of all masked sources are listed in table 1.
@@
We note that we did not mask optically selected clusters with $\lambda<20$, in which richness range the identified clusters or groups usually suffer from significantly high contamination due to projection effects (Costanzi et al. 2019; Grandis et al. 2021). Masking these sources out will significantly reduce the useful solid angle in the analysis footprint. Given that the $\lambda<20$ halos and the unmasked sources below the X-ray detection limit also contribute to the stacked signal in this work due to their associations with the large-scale structure of the Universe, we estimate the fraction of these unmasked X-ray sources and model it out from our final analysis when determining the significance of the emission from the WHIM in Sect. 4.
Table 1: Summary of masked sources.
@@ 2.2 Healpix map creation
We adopted the 0.3–1.2 keV band in the imaging analysis. This band is optimized for maximizing the WHIM emission from our selected filament sample with $z\sim 0.47$ (redshifted), assuming a $kT<1$ keV gas temperature. We split the analysis footprint into 27 tiles, each size of 10 deg $\times$ 10 deg. For each tile, we merge the event list using evtool and obtain flare-free time interval using flaregti with parameters pimin=5000, threshold=1.1, source_size=150. We use evtool and criteria PATTERN=15, FLAG=0xE [^0] F000 and GTI=FLAREGTI to select events in the flare-free time interval and generate count images with a pixel size of $16^{\prime\prime}$. We use expmap to create a corresponding vignetting corrected exposure map for each count image. After masking sources, we bin the tile image pixels to NSIDE=1024 HEALPix [^2] http://healpix.sourceforge.net (Górski et al. 2005; Zonca et al. 2019) map pixels. The NSIDE=1024 HEALPix has a $3.4^{\prime}$ pixel size corresponding to 1.4 Mpc at $z=0.6$. For the count map, the count number of the $i$ th pixel, $C_{\mathrm{HP},i}=\sum_{j}C_{ij}$, where $C_{ij}$ is the count number of the $j$ th image pixel that is located in the $i$ th HEALPix pixel. For the exposure map, the exposure time of the $i$ th pixel, $t_{\mathrm{HP},i}=(\sum_{j}t_{ij}\times\Omega_{ij})/{\Omega_{\mathrm{HP},i}}$, where $\sum_{j}t_{ij}$ and $\Omega_{ij}$ are the exposure and solid angle of the $j$ th unmasked image pixel that is located in the $i$ th HEALPix pixel, and $\Omega_{\mathrm{HP},i}$ is the solid angle of the $i$ th HEALPix pixel.
@@ 3 Surface brightness stacking
We use $i$, $j$, and $k$ to denote the indices of pixel, filament, and radial transverse distance bin, respectively, and we have $n_{\mathrm{HP}}$ HEALPix pixels, $n_{\mathrm{fil}}$ filaments and $n_{\mathrm{bin}}$ distance bins. For each combination of the $i$ th pixel $\mathcal{P}_{i}$, the $j$ th filament $\mathcal{F}_{j}$ and the $k$ th distance bin $\mathcal{B}_{k}$, we assign a Boolean coefficient $b_{ijk}$, whose value is 1 (or 0) if $\mathcal{P}_{i}$ is in (or not in) the $k$ th distance bin of $\mathcal{F}_{j}$. Therefore, for each filament $\mathcal{F}_{j}$, the count number in the $k$ th transverse distance bin is $C_{jk}=\sum_{i}^{n_{\mathrm{HP}}}{b_{ijk}\times C_{i}}$, and the exposure time in the $k$ th transverse distance bin is $t_{jk}=\sum_{i}^{n_{\mathrm{HP}}}{b_{ijk}\times t_{i}}$. We use exposure-weighted averages to obtain the stacked surface brightness profile. The stacked profile
@@
is the quotient between the sum of the count profile and the sum of the exposure profile. We use the averaged surface brightness in the outer part of the stacked profile as the local background $S_{\mathrm{X,bkg}}$ and subtract it from the entire profile to get a net surface brightness profile as
@@
Due to the projection of filaments at different redshifts, on average, each pixel is shared by 12 filaments if the profile is extracted up to 20 Mpc. It means that, for a single pixel, distance bins of different filaments at different redshifts could be simultaneously located at its position. Therefore, the surface brightness in different bins is possibly correlated. We apply a bootstrapping method to construct resampled pixel sets to obtain the estimates of the mean and covariance of the stacked profile. We use $l$ to denote the index of the randomized pixel sample. The $l$ th resample of the pixel set with replacement is
@@
Following equations 2 and 3, for each resampled pixel set $\boldsymbol{\mathcal{P}}_{l}^{{}^{\prime}}$, we obtain a stacked surface brightness profile $\boldsymbol{S}_{\mathrm{X,net},l}^{{}^{\prime}}(r_{k})$. After drawing $n_{\mathrm{bs}}$ random resamples, the bootstrapping estimate of the net surface brightness profile is
@@ 3.1 Stacking the surface brightness profile
We applied our stacking method to the HEALPix count and exposure maps created in Sect. 2.2. Fig. 3 shows the stacked surface brightness against the transverse distance from the filament spine in the 0.3–1.2 keV band. The profile is similar to those in Tanimura et al. (2020b, 2022), exhibiting a centrally peaked signal, and the intensity gradually decreases to $\sim$10 Mpc in the transverse distance. We use the 10–20 Mpc distance region to evaluate the local background level and subtract it from the profile to obtain the net surface brightness profile of the filaments. To evaluate the significance of detecting the source signal by accounting for the magnitude of fluctuations and systematic effects due to varying background levels across the sky, we construct 500 control samples of filaments. Each control sample contains 7817 filaments, but we moved them to random positions in the analysis footprint and then rotated them by random angles. The mean surface brightness profile and the corresponding standard deviation of these 500 control samples are also plotted in Fig. 3. The excess X-ray emission from the source filaments over the control sample, i.e., background in this case, is significant, reaching a $9.0\sigma$ confidence level in the 0–10 Mpc bins.
To investigate the dependence of the significance of the stacked profile on the selection of the energy band and filament lengths, we apply our analysis of two sub-bands and two distinct filament populations with different lengths and find consistent results with our analysis with default selection (see Appendix A for details).
Figure 3: Stacked filament surface brightness profile in the 0.3–1.2 keV band at the observer’s frame. The X-axis is the transverse distance from the filament spine. The orange and purple lines and shaded regions denote the profiles of the source and the averaged value of the random-positioned control samples, respectively. The average surface brightness in the 10–20 Mpc transverse distance range is used as the local background and is subtracted from the profile to obtain the surface brightness signal of the net source profile. The shaded area of the source profile is the $1\sigma$ uncertainty estimated from a pixel-wise bootstrapping method, and that of the control samples is the $1\sigma$ scatter. The correlation matrix of the 5 bins in the 0–10 Mpc range is plotted in the inset figure.
@@ 3.2 Stacking method validation
Filaments at different redshifts could be overlapped with each other due to projection effects. There are two different cases of overlapping. In the first case, the two overlapped filaments have different orientations (left in Fig. 4). When extracting the profile of filament A, the contribution from filament B in the overlapped source region will be accounted for by the filament B emission in the local background region, and vice versa. In the second case, the two overlapped filaments have similar orientations (right in Fig. 4), and the emission in the overlapping source region will be double counted and therefore increase the stacked signal. Meanwhile, our stacking method assumes that systematic uncertainties from the spatial variations of multiple components, e.g. the Milky Way foreground emission and absorption, the Cosmic X-ray background (CXB) emission, as well as the inhomogeneity of exposure is minimized by a large number of filaments randomly distributed in the footprint.
@@
To evaluate the impact of the second case overlapping and multiple possible systematic uncertainties, we first simulate mock count maps and then apply the same surface brightness profile stacking method to the mock maps to obtain the averaged stacked result and to compare it with the injected model. The details of mock map creation are presented in Appendix C. We simulate 200 mock count maps and apply the same profile stacking method to all. The stacked mock profiles from the mock maps are plotted in Fig. 5, together with the injected model profile. The averaged mock profile in general agrees with the injected model within the scatter and is slightly higher than the injected model by $13\%$. This $13\%$ enhanced signal could be due to the second case of overlapping, and we correct this effect when modeling gas properties in the subsequent sections. The validation also suggests although the averaged filament signal at the spine is $0.5\%$ of the total background level, the emission profile can be successfully recovered by our stacking method.
@@
Figure 4: Sketches of two types of overlapped filaments due to the projection effects. Deep and shallow regions are the source and local background regions, respectively. In the first case (left), the two filaments have different orientations and the emission in the overlapping region will be accounted for by local background subtraction. In the second case (right), the two filaments have similar orientations and will lead to overestimated stacked results, which is evaluated in Sect. 3.2.
@@
Figure 5: The results of 200 net surface brightness profiles of the simulated maps by applying our stacking method (in gray). The red dashed curve is the injected surface brightness profile model. The curve in purple and the shaded region are the mean and standard deviation of the 200 profiles. Although the filament signal at the spine is only $\sim 0.5\%$ of the background value, the stacking method can successfully recover the injected profile.
@@ 4 Modeling the contamination from unresolved extragalactic sources
The source masks applied in this work successfully mask out all X-ray-detected point sources and galaxy clusters and groups above the eRASS:4 limit, as well as optically selected galaxy clusters and groups with $\lambda>20$. We note that X-ray sources outside these selection criteria (i.e., X-ray sources below the eRASS:4 detection limit and low richness clusters and groups) may also contribute to the stacked X-ray signal. To remove the contaminating signal, we evaluate the contribution from three types of unmasked X-ray signal associated with galaxies: gravitationally bound hot gas in galaxy or galaxy group halos, undetected low luminous active galactic nuclei, and X-ray binaries. The steps of modeling the contribution from these sources in the stacked total signal are:
@@
    We build up a redshift dependent $L_{\mathrm{X}}-M_{*}$ scaling relation of each type of the sources in a $129.5<\log(M_{*}/M_{\odot})<$ 12 stellar mass range. We describe the contribution from each type of unmasked source in detail in Sect. 4.1.
@@
    We adopt the DESI Legacy Survey DR9 galaxy photometric redshift and stellar mass catalog (Zou et al. 2019), which is relatively complete and covers our analysis footprint. We select galaxies in the redshift range of [0.15, 0.7] and clean up this catalog by removing sources with redshift uncertainty $>20\%$. We predict the expected count rate in the 0.3–1.2 keV band for each galaxy using the constructed scaling relations, given the stellar mass, redshift, spectral shape of the source, and the Galactic absorption at the position of the galaxy. We further correct for imcompleteness of the observed stellar mass function (SMF) in Zou et al. (2019) catalog to a fiducial empirical SMF model from the result of UniverseMachine (Behroozi et al. 2019), which is a project characterizing the link between galaxy growth and halo growth. In the $M_{*}<10^{11}M_{\sun}$ range, where the galaxies in Zou et al. (2019) catalog are not complete, the ratio between the SMF of the selected galaxies from Zou et al. (2019) and the fiducial SMFs at different stellar masses and redshifts are used to correct the predicted count rate of each galaxy.
@@
    For each selected galaxy, we add the predicted count rate to the HEALPix pixel that the galaxy is located in. By this method, we create a predicted count rate map of unmasked galaxy X-ray emission in the selected redshift range. We apply the same source mask used in the profile extracting and stacking pipeline to extract the profile of unmasked signals.
Figure 6: Rest-frame galaxy averaged 0.5–2 keV luminosity - stellar mass scaling relations for unmasked X-ray halos (top left), AGNs (top right), XRBs (bottom left) at different redshifts. The scaling relations of total unmasked X-ray signals associated with galaxies are plotted in the bottom right panel, together with the recent eRASS stacking results of $z<0.2$ central galaxy sample in black.

### v2

@@ 2.1 Source masking
We masked out all eRASS:4 X-ray sources in the 0.2–2.3 keV eSASS detection catalog [^1] Proprietary eROSITA data. Version all_s4_SourceCat1B_221031_poscorr_mpe_photom. The detection pipeline used in this work is the same as the version used for the first release of the All-Sky Survey data (Merloni et al. 2024). The catalog contains point-like and extended sources, such as supernova remnants, galaxies, galaxy groups, and clusters. To remain conservative, we did not apply selection cuts to the contaminating X-ray emitting sources.
We used a background level-dependent radius to exclude detected sources as follows. For each source, the surface brightness profile can be approximately described as
@@
where $r_{\mathrm{PSF}}=9^{\prime\prime}$ is the survey averaged point spread function (PSF) core radius parameter obtained from point source stacking and $r_{\mathrm{EXT}}$ is the parameter EXT given by the catalog if the source is extended. We adopted the radius at which the source profile reaches 0.1 (10%) of the background level as the masking radius, namely, $S(r_{\mathrm{cut}})=0.1\times S_{\mathrm{bkg}}$, where $S_{\mathrm{bkg}}$ is obtained from ML_BKG_1 from the detection pipeline.
@@
For extended sources with large angular sizes, the eSASS detection pipeline usually returns several detections known as split sources (Liu et al. 2022, 2024). These large angular size sources are already included in the eRASS1 galaxy cluster and group catalog (Bulbul et al. 2024). Therefore, we masked out all sources in the eRASS1 galaxy cluster and group catalog with radii of $1.5\times r_{500}$. Additionally, we adopted an optical cluster catalog compiled by the redMaPPer algorithm (Rykoff et al. 2014) on the DESI Legacy Survey (Dey et al. 2019) DR9 + DR10 grz images in a blind mode (Kluge et al. 2024). We excluded sources with a richness of $\lambda>20$ and redshift of $0.80.2<z<$ 0.8 at radii of $1.5\times r_{\lambda}$, where $r_{\lambda}=(\lambda/100)^{1/5}$ $h^{-1}$ Mpc (this is the richness scaled radius from the redMaPPer algorithm). This cut safely removes clusters of galaxies with high completeness and high purity, especially for high-redshift systems whose X-ray luminosity is below the eRASS1 detection limit.
@@
To conservatively remove the emission from low-richness galaxy groups in the local Universe, we also adopted a galaxy group catalog (Tinker 2021) based on the SDSS Main Galaxy Sample (Strauss et al. 2002). We exclude sources with $n_{\mathrm{sat}}>1$ in radii of $1.5\times r_{500}$. If a source displays $4001.5\times r_{500}<$ 400 kpc, we used a fixed value of 400 kpc instead. In addition to those cluster and group catalogs, we masked out critical points identified by DisPerSE of type maximum and bifurcation in the filament catalog in a redshift range of $0.60.2<z<$ 0.6 with fixed radii of 2 Mpc. This range is what we consider as the nodes of the large-scale structure (see Malavasi et al. 2020, for details). Details on all masked sources are listed in table 1.
@@
We note that we did not mask optically selected clusters with $\lambda<20$, which is the richness range where the identified clusters or groups are usually seen to suffer from significantly high contamination due to projection effects (Costanzi et al. 2019; Grandis et al. 2021). Masking these sources out will significantly reduce the useful solid angle in the analysis footprint. Given that the $\lambda<20$ halos and the unmasked sources below the X-ray detection limit also contribute to the stacked signal in this work (due to their associations with the large-scale structure of the Universe), we estimated the fraction of these unmasked X-ray sources and model it out from our final analysis when determining the significance of the emission from the WHIM (see Sect. 4).
Table 1: Summary of the masked sources.
@@ 2.2 Healpix map creation
We adopted the 0.3–1.2 keV band in the imaging analysis. This band is optimized for maximizing the WHIM emission from our selected filament sample with $z\sim 0.47$ (redshifted), assuming a $kT<1$ keV gas temperature. We split the analysis footprint into 27 tiles, each size of 10 deg $\times$ 10 deg. For each tile, we merged the event list using evtool and obtain flare-free time interval using flaregti with parameters pimin=5000, threshold=1.1, and source_size=150. We used evtool and criteria PATTERN=15, FLAG=0xE [^0] F000, and GTI=FLAREGTI to select events in the flare-free time interval and generate count images with a pixel size of $16^{\prime\prime}$. We used expmap to create a corresponding vignetting corrected exposure map for each count image. After masking the sources, we binned the tile image pixels to NSIDE=1024 HEALPix [^2] http://healpix.sourceforge.net (Górski et al. 2005; Zonca et al. 2019) map pixels. The NSIDE=1024 HEALPix has a $3.4^{\prime}$ pixel size corresponding to 1.4 Mpc at $z=0.6$. For the count map, the count number of the $i$ th pixel is $C_{\mathrm{HP},i}=\sum_{j}C_{ij}$, where $C_{ij}$ is the count number of the $j$ th image pixel that is located in the $i$ th HEALPix pixel and is not masked. For the exposure map, the exposure time of the $i$ th pixel, $t_{\mathrm{HP},i}=(\sum_{j}t_{ij}\times\Omega_{ij})/{\Omega_{\mathrm{HP},i}}$, where $\sum_{j}t_{ij}$ and $\Omega_{ij}$ are the exposure and solid angle of the $j$ th image pixel that is located in the $i$ th HEALPix pixel and is not masked, and $\Omega_{\mathrm{HP},i}$ is the solid angle of the $i$ th HEALPix pixel.
@@ 3 Surface brightness stacking
We use $i$, $j$, and $k$ to denote the indices of pixel, filament, and radial transverse distance bin, respectively, and we have $n_{\mathrm{HP}}$ HEALPix pixels, $n_{\mathrm{fil}}$ filaments and $n_{\mathrm{bin}}$ distance bins. For each combination of the $i$ th pixel $\mathcal{P}_{i}$, the $j$ th filament $\mathcal{F}_{j}$ and the $k$ th distance bin $\mathcal{B}_{k}$, we assigned a Boolean coefficient $b_{ijk}$, whose value is 1 (or 0) if $\mathcal{P}_{i}$ is in (or not in) the $k$ th distance bin of $\mathcal{F}_{j}$. Therefore, for each filament, $\mathcal{F}_{j}$, the count number in the $k$th transverse distance bin is $C_{jk}=\sum_{i}^{n_{\mathrm{HP}}}{b_{ijk}\times C_{i}}$, and the exposure time in the $k$th transverse distance bin is $t_{jk}=\sum_{i}^{n_{\mathrm{HP}}}{b_{ijk}\times t_{i}}$. We used exposure-weighted averages to obtain the stacked surface brightness profile. The stacked profile,
@@
is the quotient between the sum of the count profile and the sum of the exposure profile, where $r_{k}$ is the radial transverse distance of the $k$ th bin. We used the averaged surface brightness in the outer part of the stacked profile as the local background, $S_{\mathrm{X,bkg}}$, and subtracted it from the entire profile to get a net surface brightness profile as
@@
Due to the projection effects, the distance bins of different filaments at different redshifts could be simultaneously located at a single position. Therefore, the surface brightness in different bins is possibly correlated. We applied a bootstrapping method to construct resampled pixel sets to obtain the estimates of the mean and covariance of the stacked profile. We used $l$ to denote the index of the randomized pixel sample. The $l$th resample of the pixel set with the replacement is
@@
Following Eqs. 2 and 3, for each resampled pixel set $\boldsymbol{\mathcal{P}}_{l}^{{}^{\prime}}$, we obtained a stacked surface brightness profile $\boldsymbol{S}_{\mathrm{X,net},l}^{{}^{\prime}}(r_{k})$. After drawing $n_{\mathrm{bs}}$ random resamples, the bootstrapping estimate of the net surface brightness profile is
@@ 3.1 Stacking the surface brightness profile
We applied our stacking method to the HEALPix count and exposure maps we created (as described in Sect. 2.2). Figure 3 shows the stacked surface brightness against the transverse distance from the filament spine in the 0.3–1.2 keV band. The profile is similar to those in Tanimura et al. (2020b, 2022), exhibiting a centrally peaked signal, and the intensity gradually decreases to $\sim$10 Mpc in the transverse distance. We used the 10–20 Mpc distance region to evaluate the local background level and subtracted it from the profile to obtain the net surface brightness profile of the filaments. To evaluate the significance of detecting the source signal by accounting for the magnitude of fluctuations and systematic effects due to varying background levels across the sky, we constructed 500 control samples of filaments. Each control sample contains 7817 filaments, but we moved them to random positions in the analysis footprint and then rotated them by random angles. The mean surface brightness profile and the corresponding standard deviation of these 500 control samples are also plotted in Fig. 3. The excess X-ray emission from the source filaments over the control sample (i.e.,background in this case) is significant, reaching a $9.0\sigma$ confidence level in the 0–10 Mpc bins.
To investigate the dependence of the significance of the stacked profile on the selection of the energy band and filament lengths, we applied our analysis to two sub-bands and two distinct filament populations of different lengths and found consistent results with our analysis with the default selection (see Appendix A for details). The signals in the 0.3–0.7 keV and the 0.7–1.2 keV bands are significant. Meanwhile, the detected filament width of the 20–40 Mpc lengths is marginally broader than that of the 40–100 Mpc length.
Furthermore, we investigated the impact of different source mask radii on the stacked profile (see Appendix B for details). If we increase the cluster mask radii by a factor of two, the resulting profile agrees with the original one within the uncertainties and the integrated signal within the 0–10 Mpc transverse distance bin is only reduced by 7% (see the left panel of Fig. 14). If we adopt a more conservative mask radius criteria for eRASS:4 X-ray sources, which is $S(r_{\mathrm{cut}})=0.05S_{\mathrm{bkg}}$, the resulting profile agrees with the original one within $1\sigma$ uncertainties as well and the integrated signal is enhanced by 3% (see the right panel of Fig. 14). Therefore, we conclude that increasing the source mask sizes has a trivial impact on the net profile normalization, which is in agreement with the inspections presented in Tanimura et al. (2022).
Figure 3: Stacked filament surface brightness profile in the 0.3–1.2 keV band at the observer’s frame. The X-axis is the transverse distance from the filament spine. The orange and purple lines and shaded regions denote the profiles of the source and the averaged value of the random-positioned control samples, respectively. The average surface brightness in the 10–20 Mpc transverse distance range is used as the local background and is subtracted from the profile to obtain the surface brightness signal of the net source profile. The shaded area of the source profile is the $1\sigma$ uncertainty estimated from a pixel-wise bootstrapping method, and that of the control samples is the $1\sigma$ scatter. The correlation matrix of the five bins in the 0–10 Mpc range is plotted in the inset figure.
@@ 3.2 Stacking method validation
Filaments at different redshifts could be overlapped with each other due to projection effects. The average number of projected filaments on the pixels within 4 Mpc transverse distance for our selected filament sample is 2.5. This number increases to 5.5 and 11 for pixels within transverse distances of 10 Mpc and 20 Mpc, respectively. There are two different cases of overlapping. In the first case, the two overlapped filaments have different orientations (left in Fig. 4). When extracting the profile of filament A, the contribution from filament B in the overlapped source region will be accounted for by the filament B emission in the local background region and vice versa. In the second case, the two overlapped filaments have similar orientations (right in Fig. 4), while the emission in the overlapping source region will be double counted; therefore increasing the stacked signal. Meanwhile, our stacking method assumes that systematic uncertainties from the spatial variations of multiple components, for instance, the Milky Way foreground emission and absorption, cosmic X-ray background (CXB) emission. In addition, the inhomogeneity of exposure is minimized by a large number of filaments randomly distributed in the footprint.
@@
To evaluate the impact of the second case overlapping and multiple possible systematic uncertainties, we first simulated the mock count maps and applied the same surface brightness profile stacking method to the mock maps to obtain the averaged stacked result and to compare it with the injected model. The details of the mock map creation are presented in Appendix C. We simulated 200 mock count maps and applied the same profile stacking method. The stacked mock profiles from the mock maps are plotted in Fig. 5, together with the injected model profile. The averaged mock profile is in general agreement with the injected model within the scatter; whereas, it is slightly higher than the injected model by $13\%$. This $13\%$ enhanced signal could be due to the second case of overlapping and we corrected this effect when modeling the gas properties (described in the subsequent sections). The validation also suggests that although the averaged filament signal at the spine is $0.5\%$ of the total background level, our stacking method can successfully recover the emission profile.
@@
Figure 4: Sketches of two types of overlapped filaments due to the projection effects. Deep and shallow regions are the source and local background regions, respectively. In the first case (left), the two filaments have different orientations, and the emission in the overlapping region will be accounted for by local background subtraction. In the second case (right), the two filaments have similar orientations and will lead to overestimated stacked results, evaluated in Sect. 3.2.
@@
Figure 5: Results of 200 net surface brightness profiles of the simulated maps obtained by applying our stacking method (in gray). The red dashed curve is the injected surface brightness profile model. The curve in purple and the shaded region are the mean and standard deviation of the 200 profiles. Although the filament signal at the spine is only $\sim 0.5\%$ of the background value, the stacking method can recover the injected profile successfully.
@@ 4 Modeling the contamination from unresolved extragalactic sources
The source masks applied in this work successfully mask out all X-ray-detected point sources and galaxy clusters and groups above the eRASS:4 limit, as well as optically selected galaxy clusters and groups with $\lambda>20$. We note that X-ray sources outside these selection criteria (i.e., those below the eRASS:4 detection limit and low-richness clusters and groups) may also contribute to the stacked X-ray signal. To remove the contaminating signal, we evaluated the contribution from three types of unmasked X-ray signal associated with galaxies: gravitationally bound hot gas in galaxy or galaxy group halos, undetected low-luminosity active galactic nuclei, and X-ray binaries. The steps required to model the contribution from these sources in the stacked total signal are given below.
@@
    We built a redshift-dependent $L_{\mathrm{X}}-M_{*}$ scaling relation of each type of the sources in a $129.5<\log(M_{*}/M_{\odot})<$ 12 stellar mass range. We describe the contribution from each type of unmasked source in detail in Sect. 4.1.
@@
    We adopted the DESI Legacy Survey DR9 galaxy photometric redshift and stellar mass catalog (Zou et al. 2019), which is relatively complete and covers our analysis footprint. We selected galaxies in the redshift range of [0.15, 0.7] and cleaned up this catalog by removing sources with a redshift uncertainty of $>20\%$. We predicted the expected count rate in the 0.3–1.2 keV band for each galaxy using the constructed scaling relations, given the stellar mass, redshift, spectral shape of the source, and the Galactic absorption at the galaxy’s position. We further corrected for the incompleteness of the observed stellar mass function (SMF) from Zou et al. (2019) catalog to a fiducial empirical SMF model from the result of UniverseMachine (Behroozi et al. 2019), which is a project characterizing the link between galaxy growth and halo growth. In the $M_{*}<10^{11}M_{\sun}$ range, where the galaxies in Zou et al. (2019) catalog are not complete, the ratio between the SMF of the selected galaxies from Zou et al. (2019) and the fiducial SMFs at different stellar masses and redshifts are used to correct the predicted count rate of each galaxy.
@@
    For each selected galaxy, we added the predicted count rate to the HEALPix pixel that the galaxy is located in. Thanks to this method, we were able to create a predicted count rate map of unmasked galaxy X-ray emission in the selected redshift range. We applied the same source mask used in the profile extracting and stacking pipeline to extract the profile of unmasked signals.
Figure 6: Rest-frame galaxy averaged 0.5–2 keV luminosity - stellar mass scaling relations for unmasked X-ray halos (top-left), AGNs (top-right), and XRBs (bottom-left) at different redshifts. The scaling relations of total unmasked X-ray signals associated with galaxies are plotted (bottom right) together with the recent eRASS stacking results of $z<0.2$ central galaxy sample in black.